# RealDPO: Real or Not Real, that is the Preference

### v1

@@ 1 Introduction
With the advancement in computational power and the availability of large-scale data, video generation models (Yang et al., 2024b; Guo et al., 2023; Blattmann et al., 2023; Li et al., 2024; Lin et al., 2024; Wang et al., 2024b; Xing et al., 2024; Zhang et al., 2023) have made significant progress, producing more realistic and diverse visual content. However, when it comes to generating complex motions, existing models still face considerable challenges in creating motion sequences that adhere to appear natural and smooth, and align with contextual consistency. This issue becomes especially prominent in the generation of human-centric daily activity motions. As shown in Fig. 1(b), even the results generated by the state-of-the-art DiT-based model CogVideoX-5B (Yang et al., 2024b) exhibit unnatural and unrealistic movements, failing to meet human preferences for natural, smooth, and contextually appropriate actions. This prompts us to further explore how to improve the realism and rationality of complex motion generation, particularly in the domain of human motion synthesis.
@@
A straightforward solution is to collect a set of high-quality, real-world data specifically for supervised fine-tuning (SFT). However, relying exclusively on this dataset for SFT training presents certain limitations. During optimization, the model interprets the provided data as the sole correct reference, lacking awareness of where the original model’s errors stem from. This narrow focus may result in overfitting and suboptimal performance in Fig. 2. A more effective strategy would be letting the model learn from its own mistakes. By utilizing the difference between real samples (positive data) and generative samples (negative data), we can explicitly highlight the model’s errors and guide it to correct its behavior. This approach enables the model to progressively align its outputs with the desired actions represented by the positive samples, fostering continuous improvement through self-reflection. This idea aligns perfectly with Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reinforcement learning technique used in training large language models, which leverages pair-wise win-lose data to guide the learning process.
@@
In video generation, recent studies (Liu et al., 2024; Wang et al., 2024c; Xu et al., 2024a; Yuan et al., 2024; Zhang et al., 2024a) have explored training fine-grained reward models using human-annotated preference datasets, primarily through three ways: reward-weighted regression (RWR) (Wang et al., 2024c), direct preference optimization (DPO) (Liu et al., 2024), and gradient feedback (GF) (Yuan et al., 2024). However, these methods face some critical challenges when directly applied to action-centric video generation: (1) Reward Hacking: Video reward model is susceptible to reward hacking, where during the optimization process, human evaluations indicate a decline in video quality, yet the reward model continues to assign high scores. (2) Scalability Issue: Online approaches require decoding latent to pixel space, limiting their scalability for high-resolution video generation. (3) Bias Propagation: Multi-dimensional reward models may lose the ability to assess specific key metrics due to linear combinations of evaluation criteria. As shown in Fig. 1(b), the reward model cannot provide an accurate evaluation for complex motion. These limitations highlight the need for a more robust approach tailored to complex motion video generation, motivating our extension beyond traditional DPO frameworks.
@@ 2 Related Work
Reinforce Learning in Image/Video Generation. In large language models (LLMs), reward models are commonly used in Reinforcement Learning Human Feedback (RLHF), enabling LLMs to respond more naturally and generate more coherent text. Recently, there have been a series of studies(Xu et al., 2024b; Black et al., 2023; Wallace et al., 2024; Lee et al., 2023; Liang et al., 2024; Yang et al., 2024a; Clark et al., 2023; Fan et al., 2024) in image generation that incorporate human preferences into model evaluation and model alignment training, mainly focusing on improving the aesthetic quality of images. In video generation, the exploration so far is still quite limited. Most related works(Yuan et al., 2024; Wu et al., 2024; Wang et al., 2024c; Liu et al., 2024; Zhang et al., 2024a; Liu et al., 2025) are mainly focusing on using reward models trained on human-annotated synthetic data for preference learning on model-generated data. However, these methods have some limitations. For example, training reward models may suffer from hacking issues, and multi-dimensional reward models might show reduced evaluation ability in specific domains. Additionally, relying entirely on synthetic data for preference learning could hinder the model’s potential. Therefore, we propose a novel approach that transcends the limitations of reward models by incorporating real data for preference-aligned learning.

### v2

@@ 1 Introduction
With the advancement in computational power and the availability of large-scale data, video generation models (Yang et al., 2024b; Guo et al., 2023; Blattmann et al., 2023; Li et al., 2024; Lin et al., 2024; Wang et al., 2024b; Xing et al., 2024; Zhang et al., 2023; Sun et al., 2025; Cheng et al., 2025) have made significant progress, producing more realistic and diverse visual content. However, when it comes to generating complex motions, existing models still face considerable challenges in creating motion sequences that adhere to appear natural and smooth, and align with contextual consistency. This issue becomes especially prominent in the generation of human-centric daily activity motions. As shown in Fig. 1(b), even the results generated by the state-of-the-art DiT-based model CogVideoX-5B (Yang et al., 2024b) exhibit unnatural and unrealistic movements, failing to meet human preferences for natural, smooth, and contextually appropriate actions. This prompts us to further explore how to improve the realism and rationality of complex motion generation, particularly in the domain of human motion synthesis.
@@
A straightforward solution is to collect a set of high-quality, real-world data specifically for supervised fine-tuning (SFT). However, relying exclusively on this dataset for SFT training presents certain limitations. During optimization, the model interprets the provided data as the sole correct reference, lacking awareness of where the original model’s errors stem from. This narrow focus may result in overfitting and suboptimal performance in Fig. 2. A more effective strategy would be letting the model learn from its own mistakes. By utilizing the difference between real samples (positive data) and generative samples (negative data), we can explicitly highlight the model’s errors and guide it to correct its behavior. This approach enables the model to progressively align its outputs with the desired actions represented by the positive samples, fostering continuous improvement through self-reflection. This idea aligns perfectly with Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reinforcement learning technique used in training large language models, which leverages pair-wise win-lose data to guide the learning process. Notably, Wang et al. (2025a) provides a unified theoretical perspective, showing that both SFT and DPO in LLM post-training can be interpreted as implicit reward learning within the same optimal policy–reward subspace.
@@
In video generation, recent studies (Liu et al., 2024; Wang et al., 2024c; Xu et al., 2024a; Yuan et al., 2024; Zhang et al., 2024a; Wu et al., 2025; Bahng et al., 2025; Wang et al., 2025b; Chen et al., 2025) have explored training fine-grained reward models using human-annotated preference datasets, primarily through three ways: reward-weighted regression (RWR) (Wang et al., 2024c), direct preference optimization (DPO) (Liu et al., 2024; Wu et al., 2025), and gradient feedback (GF) (Yuan et al., 2024). However, these methods face some critical challenges when directly applied to action-centric video generation: (1) Reward Hacking: Video reward model is susceptible to reward hacking, where during the optimization process, human evaluations indicate a decline in video quality, yet the reward model continues to assign high scores. (2) Scalability Issue: Online approaches require decoding latent to pixel space, limiting their scalability for high-resolution video generation. (3) Bias Propagation: Multi-dimensional reward models may lose the ability to assess specific key metrics due to linear combinations of evaluation criteria. As shown in Fig. 1(b), the reward model cannot provide an accurate evaluation for complex motion. These limitations highlight the need for a more robust approach tailored to complex motion video generation, motivating our extension beyond traditional DPO frameworks.
@@ 2 Related Work
Reinforce Learning in Image/Video Generation. In large language models (LLMs), reward models are commonly used in Reinforcement Learning Human Feedback (RLHF), enabling LLMs to respond more naturally and generate more coherent text. Recently, there have been a series of studies(Xu et al., 2024b; Black et al., 2023; Wallace et al., 2024; Lee et al., 2023; Liang et al., 2024; Yang et al., 2024a; Clark et al., 2023; Fan et al., 2024; Wang et al., 2025a) in image generation that incorporate human preferences into model evaluation and model alignment training, mainly focusing on improving the aesthetic quality of images. In video generation, the exploration so far is still quite limited. Similar to our work, RDPO (Qian et al., 2025) is also a preference-optimized method for video generation, but the two approaches differ in strategy and data handling. RealDPO adopts a pure DPO framework grounded in diffusion theory and uses real videos directly for preference data, while RDPO combines DPO with supervised fine-tuning and generates both positive and negative samples through controlled noising of real videos. Other related works(Yuan et al., 2024; Wu et al., 2024; Wang et al., 2024c; Liu et al., 2024; Zhang et al., 2024a; Liu et al., 2025; Wang et al., 2025c) are mainly focusing on using reward models trained on human-annotated synthetic data for preference learning on model-generated data. However, these methods have some limitations. For example, training reward models may suffer from hacking issues, and multi-dimensional reward models might show reduced evaluation ability in specific domains. Additionally, relying entirely on synthetic data for preference learning could hinder the model’s potential. Therefore, we propose a novel approach that transcends the limitations of reward models by incorporating real data for preference-aligned learning.